Tickets to a movie were sold at $4.00 per adult and $2.50 per student. If 240 tickets were sold for a total of $765.00, how many student tickets were sold?: Tickets to a movie were sold at $4.00 per adult and $2.50 per student. If 240 tickets were sold for a total of $765.00, how many student tickets were sold?
Answer by jojo14344(888)

Let "x" = ticket for adults
"y" = ticket for students
So x+y=240 tickets ------> eqn 1
Also
($4x)+($2.50y) = $765.00 ------> eqn 2
In eqn 1 we get, x=240-y then substitute this in eqn 2,
4(240-y) + 2.50y = 765
960-4y + 2.50y = 765
960-765=4y-2.50y
195=1.5y
y=130 tickets for students
x=240-130=110 tickets for adults
in doubt? Go back eqn 2,
$4(110) + $2.50(130) = $765
$440 + $325 = $765
$765 = $765
Thank you,
Jojo